# section[1.5em]1.5em1pc

### v2

@@ 4.   Abelian $\Gamma(\mathcal{P})$-algebras, $A$-modules
In this section, we develop the notion of a module over a divided power algebra, given by a certain abelian (or square-zero) extension, analogous to the notion of module over an operadic algebra [LodayV12]*§12.3. In the setting of usual (non-unital) $\mathbb{F}$-algebras, since any vector space can be equipped with the trivial algebra structure, the category of abelian (square-zero) $\mathbb{F}$-algebras coincide with the category of $\mathbb{F}$-vector spaces. An $A$-module then becomes a vector space equipped with an $A$-action. However, we will see that in the category of divided power algebras, abelian $\Gamma(\mathcal{P})$-algebras correspond to vector spaces equipped with additional internal operations. This section contains our candidates for the definitions of abelian $\Gamma(\mathcal{P})$-algebras and of $A$-modules. We will show in Section 5 that these are indeed the right definitions (see 5.2, 5.4).
@@
    $\beta_{x}(m+m^{\prime})$ $=\beta_{x,(n)}(m+m^{\prime})=\sum_{i+j=n}\beta_{x,(i,j)}(m,m^{\prime})=\beta_{x,(n,0)}(m,m^{\prime})+\beta_{x,(0,n)}(m,m^{\prime})$
@@
This is similar to the proof of 3.4. Denote by $n_{0}p^{0}+\dots+n_{k}p^{k}$ the mod $p$ expansion of $n$. Denote by $\underline{Q}=(\underbrace{p^{0},\dots,p^{0}}_{n_{0}},\dots,\underbrace{p^{k},\dots,p^{k}}_{n_{k}})$. Denote by $y=(\prod_{j=0}^{k}n_{j}!)^{-1}$. Using Lucas’s Theorem, one has:
@@
    $\beta_{x,(0,r_{1},r_{2},\dots,r_{s})}(a_{0},a_{1},\dots,a_{s-1},m)=\beta_{x,(r_{1},r_{2},\dots,r_{s})}(a_{1},\dots,a_{s-1},m)$.
@@
    $\beta_{x,{\underline{r}}}(\lambda a_{1},a_{2},\dots,a_{s-1},m)=\lambda^{r_{1}}\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},m)\quad\text{for all }\lambda\in\mathbb{F}$.
@@
    $\beta_{x,{\underline{r}}}(\underbrace{a_{1},\dots,a_{1}}_{q_{1}},\underbrace{a_{2},\dots,a_{2}}_{q_{2}},\dots,\underbrace{a_{s-1},\dots,a_{s-1}}_{q_{s^{\prime}}},m)=\beta_{\big{(}\sum_{\sigma\in\Sigma_{{\underline{q}}^{\prime}\rhd{\underline{r}}}/\Sigma_{\underline{r}}}\sigma\cdot x\big{)},\ {\underline{q}}^{\prime}\rhd{\underline{r}}}(a_{1},a_{2},\dots,a_{s-1},m),$
@@
    $R_{s+u}=\{n-r_{s}+k-q_{u}+\alpha k+\gamma,\alpha\in[\lambda],\gamma\in[q_{u}]\}\cup\{\lambda k+(k-q_{u})+\alpha k+\gamma,\alpha\in[\lambda^{\prime}],\gamma\in[l]\}.$
    $R_{s+u+1}=\{n-r_{s}+\lambda k+(k-l^{\prime})+\alpha k+\gamma,\alpha\in[\lambda^{\prime}],\gamma\in[l^{\prime}]\}.$
@@
We will allow the notation:
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{i-1},m,a_{i+1},\dots,a_{s}):=\beta_{\sigma_{i}^{*},{\underline{r}}^{\sigma_{i}}}(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{s},m),$
@@ 5.   Beck modules over $A$
This section is devoted to a description of Beck modules over a $\Gamma(\mathcal{P})$-algebra. Recall from Definition 2.1 that a Beck module over an object $A$ is an abelian group object over $A$; see [Beck67], [Quillen70]*§2–4, [Barr96]*§6, or [Frankland10qui]*§A for more background. The main result of this section, 5.2, is that the notion of Beck module over a $\Gamma(\mathcal{P})$-algebra $A$ corresponds to the notion of $A$-module from Definition 4.5. As a particular case, we obtain 5.4, which shows that the notion of abelian $\Gamma(\mathcal{P})$-algebra from Definition 4.1 corresponds to the notion of an abelian group object in the category of $\Gamma(\mathcal{P})$-algebras.
@@
Let ${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{pr}$${A}$ be a $\Gamma(\mathcal{P})$-algebra over $A$. Denote by $M:=\ker(p)$. The data of a Beck module structure on $B$ is equivalent to the data of a map ${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{z}$${B}$, and of a multiplication map ${B\times_{A}B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$${B}$ over $A$, satisfying the following conditions:
@@
$pr(\beta_{x,{\underline{r}}}((a_{1},m_{1}),\dots,(a_{s},m_{s})))=\beta_{x,{\underline{r}}}(pr(a_{1},m_{1}),\dots,pr(a_{s},m_{s}))=\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),$
@@
Note that, as a vector space, $B\times_{A}B\cong A\oplus M\oplus M$. One can rewrite condition (4) as the equality $\mu(a,m,0)=\mu(a,0,m)=(a,m)$. Since $\mu$ is linear,
$\mu(a,m,m^{\prime})=\mu\left((a,m,0)+(0,0,m^{\prime})\right)=\mu(a,m,0)+\mu(0,0,m^{\prime})=(a,m)+(0,m^{\prime})=(a,m+m^{\prime}),$
from which the commutativity of $\mu$ is in fact necessary.
Inspecting the $\Gamma(\mathcal{P})$-algebra structure on $B\times_{A}B$ (see previous section), one has:
$\beta_{x,{\underline{r}}}\left((a_{1},m_{1},m^{\prime}_{1}),\dots,(a_{s},m_{s},m^{\prime}_{s})\right)=(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{s},m_{1},\dots,m_{s}),\\ \nabla_{x,{\underline{r}}}(a_{1}\dots,a_{s},m^{\prime}_{1},\dots,m^{\prime}_{s})).$
On the one hand, this means that:
$\mu\left(\beta_{x,{\underline{r}}}\left((a_{1},m_{1},m^{\prime}_{1}),\dots,(a_{s},m_{s},m^{\prime}_{s})\right)\right)=\mu\bigg{(}\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{s},m_{1},\dots,m_{s}),\\ \nabla_{x,{\underline{r}}}(a_{1}\dots,a_{s},m^{\prime}_{1},\dots,m^{\prime}_{s})\bigg{)},$
Which is equal to:
$\bigg{(}\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{s},m_{1},\dots,m_{s})+\nabla_{x,{\underline{r}}}(a_{1}\dots,a_{s},m^{\prime}_{1},\dots,m^{\prime}_{s})\bigg{)}.$
On the other hand, since $\mu$ is a $\Gamma(\mathcal{P})$-algebra map, one has:
$\mu(\beta_{x,{\underline{r}}}\left((a_{1},m_{1},m^{\prime}_{1}),\dots,(a_{s},m_{s},m^{\prime}_{s})\right))$ $=\beta_{x,{\underline{r}}}\left(\mu(a_{1},m_{1},m^{\prime}_{1}),\dots,\mu(a_{s},m_{s},m^{\prime}_{s})\right)$
$=\beta_{x,{\underline{r}}}\left((a_{1},m_{1}+m^{\prime}_{1}),\dots,(a_{s},m_{s}+m^{\prime}_{s})\right)$
$=\left(\beta_{x,{\underline{r}}},\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{s},m_{1}+m^{\prime}_{1},\dots,m_{s}+m^{\prime}_{s})\right),$
@@
and that
$\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{s},0,\dots,0)=0,$ (E2)
which is also a consequence of condition (1). From relation ($\beta$3), we also obtain:
@@
$\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{i-1},0,a_{i+1},\dots,a_{s},m_{1},\dots,m_{i-1},0,m_{i+1},\dots,m_{s})=0$ (E3)
@@
$\beta_{x,{\underline{r}}}\left((a_{1},m_{1}),\ldots,(a_{s},m_{s})\right)=(\beta_{x,{\underline{r}}}(a_{1},\ldots,a_{s}),0)\\ \qquad+\sum_{(l_{i}+l^{\prime}_{i}=r_{i},l^{\prime}_{i}>0)_{i}}\bigg{(}0,\nabla_{x,(l_{1},l^{\prime}_{1},l_{2},\ldots,l_{s},l^{\prime}_{s})}(a_{1},0,a_{2},0,\ldots,a_{s},0,0,m_{1},0,m_{2},\ldots,0,m_{s})\bigg{)},$
@@
$\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),0\right)\\ +\sum_{(l_{i}+l^{\prime}_{i}=r_{i},l^{\prime}_{i}>0)_{i}}\sum_{j=1}^{s}\bigg{(}0,\nabla_{x,(l_{1},l^{\prime}_{1},l_{2},\dots,l_{s},l^{\prime}_{s})}(a_{1},0,a_{2},0,\dots,a_{s},0,0,0,\dots,0,\underbrace{m_{j}}_{(2s+2j)^{\mbox{th}}},0,\dots,0)\bigg{)}.$
@@
$\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),0\right)\\ +\sum_{j=1}^{s}\sum_{l+l^{\prime}=r_{j},l^{\prime}>0}(0,\nabla_{x,{\underline{r}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},0,a_{j+1},\dots,a_{s},0,\dots,0,\underbrace{m_{j}}_{(s+j+2)^{\mbox{th}}},0,\dots,0)).$
@@
$\beta_{y,{\underline{q}}}(b_{1},\dots,b_{s-1},m):=\nabla_{y,{\underline{q}}}(b_{1},\dots,b_{s-1},0,\dots,0,\underbrace{m}_{(2q)^{\mbox{th}}}).$ (*)
@@
$\beta_{x,{\underline{r}}}\left((a_{1},m_{1}),\dots,(a_{s},m_{s})\right)=$
$(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),0)+\sum_{j=1}^{s}\sum_{l+l^{\prime}=r_{j},l^{\prime}>0}\left(0,\beta_{\sigma_{j}^{*}x,({\underline{r}}^{\sigma_{j}})\circ_{s}(l,l^{\prime})}(a_{1},\dots,a_{j-1},a_{j+1},\dots,a_{s},a_{j},m_{j})\right).$ (**∗ ∗)
@@
$M$ is an abelian algebra:
Suppose ${\underline{r}}$ has at least two non-zero integers. For the sake of clarity we will assume that $s=2$. Then we want to prove that $\beta_{x,(r_{1},r_{2})}(m_{1},m_{2})=0$. In $B=A\oplus M$, this can be rewritten $\beta_{x,(r_{1},r_{2})}\left((0,m_{1}),(0,m_{2})\right)$, and we know that:
$\beta_{x,(r_{1},r_{2})}\left((0,m_{1}),(0,m_{2})\right)$ $=(0,\nabla_{x,{\underline{r}}}(0,0,m_{1},m_{2}))$
$=(0,\nabla_{x,{\underline{r}}}(0,0,m_{1},0)+\nabla_{x,{\underline{r}}}(0,0,0,m_{2}))=0,$
using (E1) and (E3).
$M$ is an $A$-module:
@@
Only relations ($\beta$AM4) and ($\beta$AM9) are not completely straightforward. Suppose all other relations are proven.
@@
On one hand, using the assignment (**∗ ∗5), we have:
@@
$\beta_{x,{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}((a_{1},0),\dots,(a_{s^{\prime}-1},0),\underbrace{(a_{s^{\prime}},m),\dots,(a_{s^{\prime}},m)}_{s})=\left(\beta_{\sum_{\rho\in\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}}\sigma x,{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}(a_{1},\dots,a_{s^{\prime}-1},a_{s^{\prime}}),0\right)\\ +\sum_{j=1}^{s}\sum_{k=1}^{r_{j}}\left(0,\beta_{\sum_{\Sigma_{({\underline{q}}\circ_{s^{\prime}}{\underline{r}})\circ_{s^{\prime}+j-1}(r_{j}-k,k))}}\rho\tau_{j}^{*}x,{\underline{q}}\circ_{s^{\prime}}(q_{s^{\prime}}-k,k)}(a_{1},\dots,a_{s^{\prime}},m)\right).$ (5.eq.1)
@@
$\beta_{x,{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}((a_{1},0),\dots,(a_{s^{\prime}-1},0),\underbrace{(a_{s^{\prime}},m),\dots,(a_{s^{\prime}},m)}_{s})=\\ \beta_{\sum_{\sigma\in\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}}\sigma x,{\underline{q}}}\left((a_{1},0),\dots,(a_{s^{\prime}-1},0),(a_{s^{\prime}},m)\right).$
Then, the assignment **∗ ∗5 gives:
$\beta_{x,{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}((a_{1},0),\dots,(a_{s^{\prime}-1},0),\underbrace{(a_{s^{\prime}},m),\dots,(a_{s^{\prime}},m)}_{s})=(\beta_{\sum_{\sigma\in\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}}\sigma x,{\underline{q}}}(a_{1},\dots,a_{s^{\prime}}),0)+\\ \sum_{k=1}^{q_{j}}(0,\beta_{\sum_{\sigma\in\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}}\sigma x,{\underline{q}}\circ_{s^{\prime}}(q_{s^{\prime}}-k,k)}(a_{1},\dots,a_{s^{\prime}},m)).$ (5.eq.2)
@@
Observe that a set of representative for $\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}$ is given by $\Sigma_{1}^{\times n-q_{s^{\prime}}}\times Sh({\underline{r}})$, and, if $k\leq r_{j}$ a set of representative for $\Sigma_{{\underline{q}}\circ_{s^{\prime}}(q_{s^{\prime}}-k,k)}/\Sigma_{({\underline{q}}\circ_{s^{\prime}}{\underline{r}})\circ_{s^{\prime}+j-1}(r_{j}-k,k))}$ is given by:
@@
where $\overline{F}_{{\underline{r}},k}=Sh({\underline{r}})\setminus\left(\coprod_{j\in[s]}\{\rho^{\prime}\in Sh({\underline{r}}):\rho^{\prime}(r_{1}+\dots+r_{j}-k+1)=q_{s^{\prime}}-k+1\}\right)$. Note that, if $k>r_{j}$, then $\{\rho^{\prime}\in Sh({\underline{r}}):\rho^{\prime}(r_{1}+\dots+r_{j}-k+1)=q_{s^{\prime}}-k+1\}=\emptyset$.
Remains to show that $\overline{F}_{{\underline{r}},k}=\overline{E}_{{\underline{r}},k}$. But, for $\rho^{\prime}\in Sh({\underline{r}})$, $\rho^{\prime}$ does not satisfy $\rho^{\prime}(r_{1}+\dots+r_{j}-k+1)=q_{s^{\prime}}-k+1$ for any $j\in[s]$ if and only if $(\rho^{\prime})^{-1}(\{q_{s^{\prime}}-k+1,\dots,q_{s^{\prime}}\})$ intersects at least two of the blocks $(r_{1},\dots,r_{s})$, and this happens if and only if it satisfies

### v3

@@ 4.   Abelian $\Gamma(\mathcal{P})$-algebras, $A$-modules
In this section, we develop the notion of a module over a divided power algebra, given by a certain abelian (or square-zero) extension, analogous to the notion of module over an operadic algebra [LodayV12]*§12.3. In the setting of usual (non-unital) $\mathbb{F}$-algebras, since any vector space can be equipped with the trivial algebra structure, the category of abelian (square-zero) $\mathbb{F}$-algebras coincide with the category of $\mathbb{F}$-vector spaces. An $A$-module then becomes a vector space equipped with an $A$-action. However, we will see that in the category of divided power algebras, abelian $\Gamma(\mathcal{P})$-algebras correspond to vector spaces equipped with additional internal operations. This section contains our candidates for the definitions of abelian $\Gamma(\mathcal{P})$-algebras and of $A$-modules. We will show in Section 5 that these are indeed the right definitions (see Theorem 5.2, Corollary 5.4).
@@
    $\beta_{x}(m+m^{\prime})$ $=\beta_{x,(n)}(m+m^{\prime})=\sum_{i+j=n}\beta_{x,(i,j)}(m,m^{\prime})$
    $=\beta_{x,(n,0)}(m,m^{\prime})+\beta_{x,(0,n)}(m,m^{\prime})$
@@
This is similar to the proof of Proposition 3.4. Denote by $n_{0}p^{0}+\dots+n_{k}p^{k}$ the mod $p$ expansion of $n$. Denote by $\underline{Q}$ the partition $(\underbrace{p^{0},\dots,p^{0}}_{n_{0}},\dots,\underbrace{p^{k},\dots,p^{k}}_{n_{k}})$. Denote by $y=(\prod_{j=0}^{k}n_{j}!)^{-1}$. Using Lucas’s Theorem, one has:
@@
    $\beta_{x,(0,r_{1},r_{2},\dots,r_{s})}(a_{0},a_{1},\dots,a_{s-1},m)=\beta_{x,(r_{1},r_{2},\dots,r_{s})}(a_{1},\dots,a_{s-1},m)$
@@
    $\beta_{x,{\underline{r}}}(\lambda a_{1},a_{2},\dots,a_{s-1},m)=\lambda^{r_{1}}\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s-1},m)$ for all $\lambda\in\mathbb{F}$.
@@
    $\beta_{x,{\underline{r}}}(\underbrace{a_{1},\dots,a_{1}}_{q_{1}},\underbrace{a_{2},\dots,a_{2}}_{q_{2}},\dots,\underbrace{a_{s-1},\dots,a_{s-1}}_{q_{s^{\prime}}},m)=\\ \beta_{\big{(}\sum_{\sigma\in\Sigma_{{\underline{q}}^{\prime}\rhd{\underline{r}}}/\Sigma_{\underline{r}}}\sigma\cdot x\big{)},\ {\underline{q}}^{\prime}\rhd{\underline{r}}}(a_{1},a_{2},\dots,a_{s-1},m),$
@@
    $R_{s+u}=\left\{n-r_{s}+k-q_{u}+\alpha k+\gamma,\alpha\in[\lambda],\gamma\in[q_{u}]\right\}\cup\left\{\lambda k+(k-q_{u})+\alpha k+\gamma,\alpha\in[\lambda^{\prime}],\gamma\in[l]\right\}$
    $R_{s+u+1}=\left\{n-r_{s}+\lambda k+(k-l^{\prime})+\alpha k+\gamma,\alpha\in[\lambda^{\prime}],\gamma\in[l^{\prime}]\right\},$
@@
We will use the notation:
$\beta_{x,{\underline{r}}}(a_{1},\dots,a_{i-1},m,a_{i+1},\dots,a_{s})\coloneqq\beta_{\sigma_{i}^{*},{\underline{r}}^{\sigma_{i}}}(a_{1},\dots,a_{i-1},a_{i+1},\dots,a_{s},m),$
@@ 5.   Beck modules over $A$
This section is devoted to a description of Beck modules over a $\Gamma(\mathcal{P})$-algebra. Recall from Definition 2.1 that a Beck module over an object $A$ is an abelian group object over $A$; see [Beck67], [Quillen70]*§2–4, [Barr96]*§6, or [Frankland10qui]*§A for more background. The main result of this section, Theorem 5.2, is that the notion of Beck module over a $\Gamma(\mathcal{P})$-algebra $A$ corresponds to the notion of $A$-module from Definition 4.5. As a particular case, we obtain Corollary 5.4, which shows that the notion of abelian $\Gamma(\mathcal{P})$-algebra from Definition 4.1 corresponds to the notion of an abelian group object in the category of $\Gamma(\mathcal{P})$-algebras.
@@
Let ${B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{pr}$${A}$ be a $\Gamma(\mathcal{P})$-algebra over $A$. Denote by $M\coloneqq\ker(p)$. The data of a Beck module structure on $B$ is equivalent to the data of a map ${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{z}$${B}$, and of a multiplication map ${B\times_{A}B\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mu}$${B}$ over $A$, satisfying the following conditions:
@@
$pr(\beta_{x,{\underline{r}}}((a_{1},m_{1}),\dots,(a_{s},m_{s})))$ $=\beta_{x,{\underline{r}}}(pr(a_{1},m_{1}),\dots,pr(a_{s},m_{s}))$
$=\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),$
@@
Note that, as a vector space, $B\times_{A}B\cong A\oplus M\oplus M$. One can rewrite condition (4) as the equality $\mu(a,m,0)=\mu(a,0,m)=(a,m)$. Classical computations on Beck modules for abelian groups show that, in fact, $\mu(a,m,m^{\prime})=(a,m+m^{\prime})$, from which the commutativity of $\mu$ is in fact necessary.
Inspecting the $\Gamma(\mathcal{P})$-algebra structure on $B\times_{A}B$ (see previous section), this implies that:
$\mu\left(\beta_{x,{\underline{r}}}\left((a_{1},m_{1},m^{\prime}_{1}),\dots,(a_{s},m_{s},m^{\prime}_{s})\right)\right)=$ $\bigg{(}\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{s},m_{1},\dots,m_{s})+\nabla_{x,{\underline{r}}}(a_{1}\dots,a_{s},m^{\prime}_{1},\dots,m^{\prime}_{s})\bigg{)}.$
However, since $\mu$ is a $\Gamma(\mathcal{P})$-algebra map, one also has:
$\mu(\beta_{x,{\underline{r}}}\left((a_{1},m_{1},m^{\prime}_{1}),\dots,(a_{s},m_{s},m^{\prime}_{s})\right))=\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{s},m_{1}+m^{\prime}_{1},\dots,m_{s}+m^{\prime}_{s})\right)$
@@
and that $\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{s},0,\dots,0)=0$, which is also a consequence of condition (1). From relation ($\beta$3), we also obtain:
@@
$\nabla_{x,{\underline{r}}}(a_{1},\dots,a_{i-1},0,a_{i+1},\dots,a_{s},m_{1},\dots,m_{i-1},0,m_{i+1},\dots,m_{s})=0.$ (E3)
@@
$\beta_{x,{\underline{r}}}\left((a_{1},m_{1}),\ldots,(a_{s},m_{s})\right)=\left(\beta_{x,{\underline{r}}}(a_{1},\ldots,a_{s}),0\right)+$ $\sum_{(l_{i}+l^{\prime}_{i}=r_{i},l^{\prime}_{i}>0)_{i}}\bigg{(}0,\nabla_{x,(l_{1},l^{\prime}_{1},l_{2},\ldots,l_{s},l^{\prime}_{s})}(a_{1},0,a_{2},0,\ldots,a_{s},0,0,m_{1},0,m_{2},\ldots,0,m_{s})\bigg{)},$
@@
$\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),0\right)+$ $\sum_{(l_{i}+l^{\prime}_{i}=r_{i},l^{\prime}_{i}>0)_{i}}\sum_{j=1}^{s}\bigg{(}0,\nabla_{x,(l_{1},l^{\prime}_{1},l_{2},\dots,l_{s},l^{\prime}_{s})}(a_{1},0,a_{2},0,\dots,a_{s},0,0,0,\dots,0,\underbrace{m_{j}}_{(2s+2j)^{\mbox{th}}},0,\dots,0)\bigg{)}.$
@@
$\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),0\right)+$ $\sum_{j=1}^{s}\sum_{l+l^{\prime}=r_{j},l^{\prime}>0}(0,\nabla_{x,{\underline{r}}\circ_{j}(l,l^{\prime})}(a_{1},\dots,a_{j},0,a_{j+1},\dots,a_{s},0,\dots,0,\underbrace{m_{j}}_{(s+j+2)^{\mbox{th}}},0,\dots,0)).$
@@
$\beta_{y,{\underline{q}}}(b_{1},\dots,b_{s-1},m)\coloneqq\nabla_{y,{\underline{q}}}(b_{1},\dots,b_{s-1},0,\dots,0,\underbrace{m}_{(2q)^{\mbox{th}}}).$ (*)
@@
$\beta_{x,{\underline{r}}}\left((a_{1},m_{1}),\dots,(a_{s},m_{s})\right)=\left(\beta_{x,{\underline{r}}}(a_{1},\dots,a_{s}),0\right)+\\ \sum_{j=1}^{s}\sum_{l+l^{\prime}=r_{j},l^{\prime}>0}\left(0,\beta_{\sigma_{j}^{*}x,({\underline{r}}^{\sigma_{j}})\circ_{s}(l,l^{\prime})}(a_{1},\dots,a_{j-1},a_{j+1},\dots,a_{s},a_{j},m_{j})\right).$ (5.eq.1)
@@
Let us show that $M$ is an abelian algebra:
Suppose ${\underline{r}}$ has at least two non-zero integers. For the sake of clarity we will assume that $s=2$. Then we want to prove that $\beta_{x,(r_{1},r_{2})}(m_{1},m_{2})=0$. In $B=A\oplus M$, this can be rewritten $\beta_{x,(r_{1},r_{2})}\left((0,m_{1}),(0,m_{2})\right)$, which, using (E1) and (E3), is equal to 00.
Let us show that $M$ is an $A$-module:
@@
Only relations ($\beta$AM4) and ($\beta$AM9) are not straightforward.
@@
On one hand, using the assignment (5.eq.1), we have:
@@
$\beta_{x,{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}((a_{1},0),\dots,(a_{s^{\prime}-1},0),\underbrace{(a_{s^{\prime}},m),\dots,(a_{s^{\prime}},m)}_{s})=\left(\beta_{\sum_{\rho\in\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}}\sigma x,{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}(a_{1},\dots,a_{s^{\prime}-1},a_{s^{\prime}}),0\right)+\\ \sum_{j=1}^{s}\sum_{k=1}^{r_{j}}\left(0,\beta_{\sum_{\Sigma_{({\underline{q}}\circ_{s^{\prime}}{\underline{r}})\circ_{s^{\prime}+j-1}(r_{j}-k,k))}}\rho\tau_{j}^{*}x,{\underline{q}}\circ_{s^{\prime}}(q_{s^{\prime}}-k,k)}(a_{1},\dots,a_{s^{\prime}},m)\right).$ (5.eq.2)
@@
$\beta_{x,{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}((a_{1},0),\dots,(a_{s^{\prime}-1},0),\underbrace{(a_{s^{\prime}},m),\dots,(a_{s^{\prime}},m)}_{s})=\beta_{\sum_{\sigma\in\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}}\sigma x,{\underline{q}}}\left((a_{1},0),\dots,(a_{s^{\prime}-1},0),(a_{s^{\prime}},m)\right).$
Then, the assignment (5.eq.1) gives:
$\beta_{x,{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}((a_{1},0),\dots,(a_{s^{\prime}-1},0),\underbrace{(a_{s^{\prime}},m),\dots,(a_{s^{\prime}},m)}_{s})=\\ (\beta_{\sum_{\sigma\in\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}}\sigma x,{\underline{q}}}(a_{1},\dots,a_{s^{\prime}}),0)+\sum_{k=1}^{q_{j}}(0,\beta_{\sum_{\sigma\in\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}}\sigma x,{\underline{q}}\circ_{s^{\prime}}(q_{s^{\prime}}-k,k)}(a_{1},\dots,a_{s^{\prime}},m)).$ (5.eq.3)
@@
Observe that a set of representative for $\Sigma_{{\underline{q}}}/\Sigma_{{\underline{q}}\circ_{s^{\prime}}{\underline{r}}}$ is given by $\Sigma_{1}^{\times n-q_{s^{\prime}}}\times Sh({\underline{r}})$, and, if $k\leq r_{j}$ a set of representative for
$\Sigma_{{\underline{q}}\circ_{s^{\prime}}(q_{s^{\prime}}-k,k)}/\Sigma_{({\underline{q}}\circ_{s^{\prime}}{\underline{r}})\circ_{s^{\prime}+j-1}(r_{j}-k,k))}$
is given by:
@@
where
$\overline{F}_{{\underline{r}},k}=Sh({\underline{r}})\setminus\left(\coprod_{j\in[s]}\{\rho^{\prime}\in Sh({\underline{r}}):\rho^{\prime}(r_{1}+\dots+r_{j}-k+1)=q_{s^{\prime}}-k+1\}\right).$
Note that, if $k>r_{j}$, then $\{\rho^{\prime}\in Sh({\underline{r}}):\rho^{\prime}(r_{1}+\dots+r_{j}-k+1)=q_{s^{\prime}}-k+1\}=\emptyset$.
It remains to show that $\overline{F}_{{\underline{r}},k}=\overline{E}_{{\underline{r}},k}$. But, for all $\rho^{\prime}\in Sh({\underline{r}})$, the condition for $\rho^{\prime}$ not to satisfy $\rho^{\prime}(r_{1}+\dots+r_{j}-k+1)=q_{s^{\prime}}-k+1$ for any $j\in[s]$ is equivalent to asking that $(\rho^{\prime})^{-1}(\{q_{s^{\prime}}-k+1,\dots,q_{s^{\prime}}\})$ intersects at least two of the blocks $(r_{1},\dots,r_{s})$, and this happens if and only if $\rho^{\prime}$ satisfies